### v2

# StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models

Abstract
Planning and acting to solve ‘real’ tasks using large language models (LLMs) in interactive environments has become a new frontier for AI methods. While recent advances allowed LLMs to interact with online tools and solve robotics tasks, long range reasoning tasks remain a problem for LLMs. Existing methods to address this issue are very resource-intensive and require additional data or human crafted rules; instead, we propose StateAct, a simple method based on few-shot in-context learning alone. StateAct enhances ‘chain-of-thought’ with state- and goal-tracking to help with planning and acting with LLMs. The method establishes the new state-of-the-art on Alfworld for in-context learning methods (+14% over the previous best few-shot in-context learning method) and performs on par with methods that use additional training data and additional tools such as code-execution. We also demonstrate that StateAct with enhanced ‘chain-of-states’ allows the agent to both solve longer horizon problems and to be more efficient in the number of steps required to solve a task. We show that StateAct works across a variety of API-based and open source LLMs. Finally, we also conduct ablation studies and show that ‘chain-of-thoughts’ help state-tracking accuracy, while a json-structure harms overall performance. We open-source our code and annotations at https://github.com/ai-nikolai/stateact.
StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models
Nikolai Rozanov Imperial College London Department of Computing nikolai.rozanov13@imperial.ac.uk                        Marek Rei Imperial College London Department of Computing marek.rei@imperial.ac.uk
1 Introduction
Using the in-built world and commonsense knowledge [^1] Commonsense- and world- knowledge as explored by Lauscher et al. (2020), for example. of large language models (LLMs), such as GPT-3, Mixtral and Gemini Brown et al. (2020); Jiang et al. (2024); Anil et al. (2023) to perform interactive reasoning tasks has become a frontier in AI research, with “AI Agents” now able to solve a range of multi-modal complex tasks Durante et al. (2024). These range from solving simulated robotics tasks Puig et al. (2018); Shridhar et al. (2021) and digital tasks, such as online shopping Yao et al. (2023a) and navigating operating systems Liu et al. (2023), to playing a variety of games Côté et al. (2019); Liu et al. (2023).
While LLMs are increasingly successful in solving AI agent tasks, Li et al. (2023b) show that LLMs struggle with long range understanding and Coelho et al. (2024) show that LLMs often focus on the beginning of the textual history, further amplifying the problem. Existing efforts to address these issues are resource-intensive: Wu et al. (2024) required human expert annotations of rules, Sun et al. (2023) required a code execution environment with carefully crafted code-based prompts, and Fu et al. (2024) used additional training data together with retrieval augmented generation (RAG) to help the AI agent.
To overcome these challenges, we introduce “StateAct”, a novel LLM-agent based on few-shot in-context learning. The method tracks the goal by ‘reminding’ the agent of the goal. StateAct, also, explicitly keeps track of the agent’s state (such as location and inventory). Additionally, we utilise ‘chain-of-thought’ (i.e. explicit reasoning Wei et al. (2023)) as an optional addition. At the core of the method lies a straightforward annotation of the few-shot examples, specifically with goal-, state- and reasoning- traces, which we call ‘chain-of-states’.
Our method establishes a new state-of-the-art for Alfworld Shridhar et al. (2021) for few-shot in-context learning-based methods, outperforming the previous best method by 14% and even outperforming methods that use external tools by 2.5%. Furthermore, while our method does not require additional data or external tools and only minimal and easy human annotations, we perform on par with the current state-of-the-art that is much more resource-intensive and requires additional data and expert human annotations. Comparing to in-context learning methods, we also demonstrate that our method is more efficient in terms of the number of steps taken to solve a task and is able to solve longer-horizon tasks, validating the hypothesis that explicit state-tracking and goal-tracking help with solving longer horizon reasoning and planning tasks.
2 Background
AI agents have historically used reinforcement learning (RL) to solve tasks Sutton and Barto (2018). With the capabilities of LLMs improving, recent work has combined LLMs and RL, training additional policies or value functions to make predictions (Li et al., 2022; Nottingham et al., 2023).
2.1 In-context learning approaches
Huang et al. (2022a, b) and Singh et al. (2022) were among the first to use LLMs directly to act in an interactive environment; their method produces agent actions as output after receiving environment observations as input.
ReAct Yao et al. (2023b) took this work further by combining ‘acting‘ Huang et al. (2022a) and ‘chain-of-thought‘ Wei et al. (2023). ReAct establishes state-of-the-art for for approaches based only on in-context learning and while it is a very scalable method, the performance (i.e. success rate) is still limited.
ExpeL Zhao et al. (2023) uses additional training data to generate ‘insights’ and ‘success trajectories’ during training. At inference time, they look up the closest ‘success trajectories’ as few-shot examples to the agent (as opposed to the fixed few-shot examples that we use) and augment them with these ‘insights’ to perform the final inference. They achieve 59% on Alfworld using retrieved ‘success trajectories’ (as few-shot examples) + ‘insights’ and 50% using the same few-shot examples as ReAct + ‘insights’.
2.1.1 In-context learning and additional tools
The current state-of-the-art for in-context learning-based approaches in combination with additional tools is ProgPrompt (Singh et al., 2022) and follow-on work AdaPlanner Sun et al. (2023). AdaPlanner introduces a code-based prompt Li et al. (2023a). They use code-execution as an additional tool, by executing LLM generated code and feeding the results into the next LLM generation. The shortcoming of AdaPlanner is that it requires very complex human-crafted prompts that are hard to scale to new environments and it requires the additional step of code-execution.
2.2 State tracking in LLM-based agents
Concurrent work to ours, AutoGuide Fu et al. (2024), uses ReAct as the base agent as well as additional training data to create ‘state-aware’ text-based guidelines for the LLM-agent; they then use a type of retrieval augmented generation (RAG) process to guide the decision making process. They embed the current observation as a summary (which they call ‘state-aware’) and use an LLM to ‘look up’ the relevant ‘state-aware’ guideline, which is then fed into a final LLM to generate an action. Using training data and LLM-based RAG, they achieve 79.1% on top of a ReAct agent on Alfworld. Their training and RAG approach could be used complementary to our StateAct LLM-agent [^2] However, the cost of using RAG at every generation step could be significantly higher than our method..
Chen et al. (2024) propose state-tracking as a way to help the agent solve the task without training data. Their method differs to ours two-fold. Firstly, they employ a complex sequence of components working together (an LLM-based attention over the observations, an LLM-based compilation of a complex state and, finally, a prediction of a program). Secondly, their system involves execution of actual programs. Our method, on the other hand, requires a straightforward extension of ‘chain-of-thought’ and uses a single LLM call to produce the state, thought and action. Additionally, we do not require program execution.
2.2.1 Fine-tuning approaches
Previous fine-tuning approaches did not significantly enhance performance Zhou et al. (2024); Yao et al. (2023b); Shridhar et al. (2021). A concurrent work, however, ActRe Yang et al. (2024b) achieves 83% by fine-tuning gpt-3.5-1106 on additional training data.
2.3 “Multi-Agent Conversation” approaches
A new trend is to use multiple LLMs concurrently to ‘chat’ to one another to produce a result. A recent work in this direction by Wu et al. (2023) achieves 67% on Alfworld.
2.4 Joined rule and LLM-based agents
StateFlow Wu et al. (2024) uses Finite State Machines (FSMs) combined with LLMs to solve Alfworld. These FSMs are human expert-crafted states, transitions and rule-based heuristics, where the LLM is asked to perform limited tasks in each of the given states. While their performance is 82% on Alfworld, we believe this result is more comparable to a rule-based ‘expert’ than an LLM-based agent (notably close to 100% can be achieved on Alfworld using hand-crafted ‘experts’ Shridhar et al. (2021)).
2.5 Approaches with multiple trials
Works such as Reflexion Shinn et al. (2023) or work by Sun et al. (2023); Yang et al. (2024a) are methods that perform several evaluations on the same test environments and perform learning loops (during evaluation) on these test-set environments to improve the score. While the score necessarily goes up (as only the failed environments are re-run), we do not dive into this as it does not compare to the actual agent performance; additionally, ‘reflection‘ can be used on top of most LLM-based agents (e.g. Reflexion uses ReAct as the base agent).
3 Method
Figure 1: StateAct interacts with the environment directly, without additional tools or resources (a). A skeleton of the StateAct prompt (b).
StateAct is an LLM-based AI agent that works on top of pre-trained large language models. StateAct takes the textual ‘observation’ from the environment and, after a single call to the pre-trained LLM, returns the ‘action’ back to the environment (without the use of additional tools or resources), see Figure 1(a).
StateAct utilises in-context learning Brown et al. (2020); Wei et al. (2023) to make the agent interact with the environment. At the core of the approach is a prompt that consists of few-shot examples of successful ‘interaction traces’ as well as the current interaction trace up to the current step in the environment, see Figure 1(b). An interaction trace consists of alternating observations from the environment and desired (or actual) outputs from the LLM. In the case of StateAct, the LLM is tasked to generate the ‘goal’, ‘state’, ‘thought’ and ‘action’. The action is then extracted and passed to the environment to produce the next observation, see Figure 2. For example, a sample of the trace for step n in Alfworld is:
Observation n:
On the countertop 1, you see a tomato 1.
LLM Output n:
>Goal: Put a clean tomato in fridge
Current Location: countertop 1
Current Inventory: None
Thought: I find the tomato, now I need to take it.
Action: take tomato 1
In the above example, the agent produces the action ‘take tomato 1’ in response to the observation ‘On the countertop 1, you see a tomato 1’.
Figure 2: An example of how StateAct is working. The depicted traces are from the Alfworld environment.
3.1 Goal-reminders and state-tracking using LLMs
In order to make StateAct more precise, we introduce goal and state tracking. We denote $\pi$ as the policy of an AI-agent, in the standard case at time step $t$, the policy predicts action $a_{t}$, given the history of observations and actions $[o_{t},a_{t-1},...,a_{0},o_{0}]$.
$\pi(a_{t}|o_{t},a_{t-1},...,a_{0},o_{0})$ (1)
Where $a_{t}$ is the action produced by the agent at time step t and $o_{t}$ is the observation produced by the environment at time step t after receiving action $a_{t}$ as input. Usually, the first observation $o_{0}$ also contains the ‘goal’ description for the given environment.
For our case, we need to enhance the policy to incorporate the ‘state’. Similar to previous work Yao et al. (2023b), we introduce the context vector, $c_{t}$. The context vector contains the action, as well as the additional predictions of the agent , i.e. $c_{t}=(g_{0},s_{t},r_{t},a_{t})$. Where $g_{0}$ is the goal and it always remains the same (for a given environment) and uses the goal extracted from $o_{0}$; $s_{t}$ represents the predicted state at time step $t$; $r_{t}$ represents ‘chain-of-thought’-style ‘reasoning’ at time step $t$; and $a_{t}$ represents the action at time step $t$, as before. The new policy $\pi$ then becomes:
$\pi_{contextual}(c_{t}|o_{t},c_{t-1},...,c_{0},o_{0})$ (2)
In our case, the LLM acts as $\pi_{contextual}$ and produces the context vector at every time step.
4 Experimental setup
Our aim is to study the long-range acting, planning and reasoning capabilities of LLM-based agents. To achieve this, in line with previous work, we turn to simulated environments as an evaluation framework and to API-based state-of-the-art large language models. Specifically, we use Alfworld Shridhar et al. (2021), a household robotics environment, and Webshop Yao et al. (2023a), an online shopping environment, as our simulated environments. In line with previous work, we use OpenAI’s [^3] https://openai.com, last accessed June 2024. gpt-3.5-turbo-1106. We, also, evaluate Alfworld on additional LLMs and validate our results.
4.1 Alfworld
Figure 3: An example textual interaction in Alfworld (right) and corresponding 3D rendering (left).
Alfworld Shridhar et al. (2021) is based on a 3D, visual, household robotics environment called Alfred Shridhar et al. (2020), which was translated into a text-based environment for ease of use for language-based AI models, see Figure 3. Alfworld has a total of 134 test-set examples and six environment types. It features long-time horizons, partial observability, an out-of-distribution evaluation set and text-based interactions. Alfworld simulates a household environment with a household assistant robot tasked with solving problems, e.g. clean an apple and put it on a table. The robot (or agent) then needs to perform a series of ‘high-level’ operations to accomplish the tasks, e.g. ‘go to fridge 1’, ‘open fridge 1’. At every step, the environment provides either a textual observation or feedback that the command has failed, e.g. ‘You open the fridge 1’, ‘You see apple 1’. The underlying text engine is based on Textworld Côté et al. (2019), see Appendix A for a complete list of the commands and further details on the environments.
4.1.1 Alfworld correction
In our research, we identified that Alfworld has a specific syntactic feature for the put command, namely put <object> in/on <place>, where “in/on” needs to be written exactly this way. Using only “in” or only “on” produces a failed command. We observed this issue with LLMs in this environment and we propose a simple fix for it. We map: 1. “put <object> in <place>” and 2. “put <object> on <place>” to the command accepted by Alfworld, namely “put <object> in/on <place>”.
Methods such as AdaPlanner Sun et al. (2023) have avoided this issue because they use code-based prompts and regex parsers. However, methods such as ReAct Yao et al. (2023b) and ExpeL Zhao et al. (2023) have been affected by it therefore lowering their potential performance. We report results for ReAct and StateAct both with and without these corrections.
4.2 Webshop
Figure 4: An textual interaction in Webshop (right) and correspondingwebsite rendering (left).
Webshop Yao et al. (2023a) is a simulation of an online shopping experience. Given a task, e.g. “I want a blue water-proof winter jacket, less than \$100”, the agent needs to search a product catalogue, browse through the search results, select the most fitting product, select the attributes, e.g. colour, size, and then buy the product. In line with previous work, we use the text-based version of Webshop, where all descriptions of the website are given in text form, see Figure 4. Webshop features a realistic large-scale product catalogue, a search engine and very varied product attributes depending on the category of product, see Appendix B for more details. In total, the test set consists of 500 examples and each one is of the type “search and buy a product”. Overall, Webshop has a maximum of 15 steps and two commands: 1. search[<query>], 2. click[<button>].
4.3 In context learning
Since ReAct Yao et al. (2023b) forms the underlying agent for many current Zhao et al. (2023) and state-of-the-art approaches Fu et al. (2024), we use the same few-shot ‘interaction traces’ as ReAct. The main reason is to have a fair comparison and to isolate additional effects - such as performance change - from different in-context examples.
In total, Alfworld has six types of tasks and ReAct uses two in-context examples per task type to prompt the language models. On average, each ReAct example ranges from 352 words to 591 words (590 tokens to 935 tokens). For our study, we reuse the observations, thoughts and actions, and annotate these examples further with goal ‘reminders’ and state tracking, which results in a range from 484 to 911 words (807 tokens to 1458 tokens) per example. During our annotation, we discovered minor errors in the ReAct prompts and fixed them as well. We release all our annotations with our code release. In comparison, AdaPlanner Sun et al. (2023), uses a different code-based approach and the prompt has 1104 words (2015 tokens) on average.
We use the two-shot examples from ReAct for Alfworld and the one-shot example from ReAct for Webshop in all our experiments.
4.4 Models
In line with previous work, we focus our main experiments are based on the API-based LLMs to compare performance fairly. Concretely, ‘gpt-3.5’ level models were used in previous work including current state-of-the-art ActRe (Yang et al., 2024b) and state-of-the-art for in-context learning ReAct (Yao et al., 2023b). Many OpenAI models have become deprecated; notably, all models from ReAct and AdaPlanner Sun et al. (2023) davinci-002, gpt-3.5-turbo-0301, and gpt-3.5-turbo-0613 are now deprecated. Therefore, we re-implemented ReAct and ran the experiments to determine the most suitable model, see Table 1. We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this model for our main experiments.
In our work, we also compare our method using newer state-of-the-art models to show that our method generalises to new models. Specificially, we use the gpt-4o-mini [^4] Again, we chose the smaller model due to cost. and Mixtral (Jiang et al., 2024) models. We note that full GPT-4 level models are prohibitively expensive [^5] A single evaluation run on Alfworld costs approx. \$8 using gpt-3.5 and ReAct, whilst gpt-4 costs 10+ times more.. We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C for the exact settings.
Method Version AW WS
ReAct gpt-3.5-turbo-1106 40.7 18.2
ReAct gpt-3.5-turbo-0125 30.37 14.6
Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS), using available gpt-3.5 models.
4.5 Metrics
In terms of metrics, we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR). Success is a binary metric per each environment in the respective test sets (134 for Alfworld and 500 for Webshop respectively). Success in Alfworld means the agent has successfully completed the whole task. In Webshop, it means the agent has bought an item that has a one hundred percent match with the desired item based on a partially hidden list of shopping item attributes (e.g. the colour, size, price, etc.).
5 Results
5.1 Alfworld
For Alfworld, we present the results for ReAct, AdaPlanner with and without code execution, and StateAct (ours), which consists of goal + state + thought + action. We also show StateAct without each of the components (i.e. without goal, state and thought). Interestingly, we find that, contrary to previous findings, ‘thought’ or ‘reasoning’ actually sometimes hinders the performance.
In Table 2, we can see that StateAct with all goal + state + thought and the correction performs the best. It outperforms ReAct with correction by around 13 points (when using the same GPT model for ReAct) and by around 9 points (when using the better model for ReAct). StateAct also outperforms ReAct by 22 points when corrections are not used. Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command-mapping (similar to our correction) along with using code-execution.
Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16- and 23-point jump for ReAct and a 27-point jump for StateAct. This indicates that the model generally performs very well; however, it struggles with minute differences in domain-specific syntax.
Method GPT-3.5 SR %
Baselines w/o corrections
ReAct 0301* 51.9
ReAct 1106 40.7
ReAct 0125 30.37
ReAct (joined***) 1106 36.30
Results w/o corrections
StateAct (ours) 1106 50.37
- w/o thought 1106 62.96
- w/o goal 1106 42.96
- w/o state 1106 44.44
Baselines with corrections
ReAct + corr. 0301* 68.15
ReAct + corr. 1106 63.70
AdaPlanner No-Exec** 0301* 46.66
AdaPlanner 0301* 75.56
Results with corrections
StateAct + corr. (ours) 1106 77.04
- w/o thought 1106 64.44
- w/o goal 1106 74.04
- w/o state 1106 64.44
Table 2: Success Rate (SR) on the 134 test-set examples from Alfworld. *gpt-3.5-0301 was deprecated in June 2024. **No-Exec means AdaPlanner without code execution. ***joined means that thought + action are produced at every turn. AdaPlanner results are from Sun et al. (2023). All other results are single run.
5.2 Webshop
For Webshop, we present the results for ReAct and StateAct. Similarly to Alfworld, we also present the results of StateAct without each of goal, state and thought, see Table 3. Interestingly, we see that removing thought produces the highest results and outperforms ReAct by 10 points. Our hypothesis is that domain-specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
Method GPT-3.5 Success Rate %
Baselines
Rule-based* N/A 9.60
RL* N/A 17.60
ReAct 1106 17.80
Results
StateAct (ours) 1106 17.00
- w/o thought 1106 27.80
- w/o goal 1106 20.40
- w/o state 1106 21.00
Table 3: Success Rate (SR) on the 500 test-set examples from Webshop. *results taken from Yao et al. (2023a). For StateAct without thought and ReAct Results are the average of two runs. Other Results are single run.
5.3 Results using additional models
We validate our results further with additional LLMs. Again we see a significant boost in performance, see Table 4. We experimented with gpt-4o-mini and Mixtral [^6] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’. (Jiang et al., 2024) on Alfworld. ReAct achieves 68.15 and 72.59 respectively, while StateAct achieves 71.85(+3.7) and 83.70(+11.2).
Method Model SR %
Baselines
ReAct Gpt-4o-mini 68.15
ReAct Mixtral-8x22B* 72.59
Results
StateAct (ours) Gpt-4o-mini 71.85
StateAct (ours) Mixtral-8x22B* 83.70
Table 4: Success Rate (SR) on the 134 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtral is ‘mixtral-8x22b-instruct-v0.1’.
5.4 Summary of results
In conclusion, we found that our simple goal-reminding and state-tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more than 10 points and even outperforms leading approaches that rely on code-execution. Interestingly, we found that the approaches are quite sensitive to domain-specific syntax and, when this is the case, ‘thoughts’ that are verbose can hinder performance.
6 Analysis and Ablations
In the results section, we discovered that our methods perform better than previous state-of-the-art. Our method shows strong performance with in-context learning without resorting to additional tools, data or bigger models. In this section, we analyse our results further and also answer our second hypothesis that goal ‘reminding’ and state-tracking help with long-range reasoning. For all ablation studies, we focus on Alfworld as it has two favourable properties over Webshop. Firstly, Alfworld has a longer time horizon (50 steps vs. 15 in Webshop), with tasks taking, on average, less than 10 steps in Webshop and around 20 to 30 steps in Alfworld. Secondly, Alfworld has much less domain-specific syntax and is purely text-based, while Webshop has a more specific syntax to follow.
6.1 Do goal reminders help with long-range tasks?
For this purpose, we compare the original ReAct (thought + action) with the goal included, i.e. StateAct (goal + thought + action). In Figure 5 we can see that, while the performance of both ReAct and StateAct goes down as there are more steps, the goal tracking has a substantially better relative performance as the number of steps increases. This finding is in-line with our original motivation that LLM agents deteriorate in performance as the prompts and interactions get longer.
To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct [^7] We ignore ‘thought’ turns for ReAct as otherwise ReAct would have even more steps. and StateAct. Table 5 clearly show that ReAct, with an average of 38.84 steps to solve an environment, is the least efficient whilst StateAct, with an average of 28.96 steps to solve an environment, is the most efficient. This shows that not only does goal reminders help with longer range tasks, it also helps with efficiency, by shortening the tasks. See Appendix E for more discussion.
Figure 5: Goal vs. No Goal, on the test set of Alfworld, using gpt-3.5-turbo-1106 without correction.
6.2 What effects does state-tracking have?
We also analyse whether state-tracking helps with long-range reasoning and efficiency. We compare the full StateAct against StateAct without state-tracking, as well as comparing ReAct (thought + action) against StateAct with state-tracking (state + thought + action). In Figure 6 we see that state-tracking also helps with long-range reasoning. In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with both state tracking and goal-tracking longer-range tasks can be solved. Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^8] In terms of cost, we found that, despite our method using a twice-longer prompt, our cost remains similar, at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer steps.. Concretely, ReAct has 24/134 examples that are in the bucket ‘40-50’ and solves 0, while our method has 6/134 examples and solves 4.
Therefore, we find that explicit state-tracking helps with long-range tasks and enables the agent to solve the tasks more efficiently.
Figure 6: State vs. No State, on the 134 test examples from Alfworld. gpt-3.5-turbo-1106 with correction.
Method Avg. Steps $\downarrow$
ReAct 38.84
StateAct (goal+thought+action) 31.19
StateAct (full) 28.96
ReAct with corr. 31.49
StateAct with corr. 19.11
- w/o thought 23.76
- w/o goal 20.09
Table 5: Average number of Steps (Avg. Steps) [lower is better] on test-set of Alfworld, using gpt-3.5-1106.
6.3 Does the model perform actual state-tracking?
We ask ourselves whether the model is actually performing state-tracking. For that purpose, we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^9] See Appendix D for details. based on the actions the agent takes. For example, if the agent produces the action go to fridge 1 and the environment accepts this action, we update the state with current location: fridge 1. We compare the ‘gold’ state against the predicted state. Figure 7 shows that StateAct does in fact correct state-tracking 88% of the time. We also observe that thoughts and goals help the state-tracking.
Figure 7: State-tracking accuracy for StateAct on 134 test examples of Alfworld, using gpt-3.5-1106.
6.4 Does json structure help with performance?
Since we found that domain-specific syntax hinders performance, we investigated whether adding a structured format like json would help. For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix G for more details. Surprisingly, we found that the json format hinders performance significantly, see Table 6. However, we also see that corrections help the json format less, indicating that json helps with syntax but hinders performance.
Method SRSR (+json)%
StateAct 50.37 45.19 (-5.2)
StateAct w/ corr. 77.04 58.52(-18.5)
Table 6: No-json vs. json. Success Rate (SR) on the test set of Alfworld, using gpt-3.5-1106.
7 Conclusion and future work
Inspired by the challenge that LLMs and LLM agents struggle with long context, we propose a novel in context learning method for LLM agents - StateAct.
We introduce ‘chain-of-states’, which relies on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution. StateAct outperforms the previous state-of-the-art that uses in-context learning alone, by between 9% and 20% given different models and tasks, and outperforms in-context learning with tools (i.e. code-execution) by 3%, while performing on par with methods that require additional training data.
In conclusion, we show that explicit state-tracking and goal reminders help the model solve long range reasoning tasks and interestingly also make the model more efficient.
8 Ethical Considerations
8.1 Computational footprint
Running many of the experiments presented in this paper can have a significant computational footprint. We should consider the environment and financial resources for reproduciblity of our work. We aimed to address this concern by using gpt-3.5-turbo level models, reporting costs and minimising the cost of our method.
8.2 Hallucinations in LLMs
As LLM-based agents become more powerful and therefore more pervasive in our daily lives, ‘hallucinations’ in LLMs can be very harmful Wei et al. (2024). We hope that explicit state-tracking presented in this work can also lead to future work to reduce ‘hallucinations.’
9 Limitations
9.1 Languages and evaluation benchmarks
We evaluated our method only in the English language and on two evaluation benchmarks. While we do not expect major changes in other languages, this is something that should be investigated. Furthermore, performance on other benchmarks should be evaluated as well.
9.2 Reasoning traces that rely on human judgement
Our prompts require human annotations; as such, there is a natural bias present. This can have both task-performance implications as well as ethical implications.
Appendix A Alfworld
A.1 Environment Types
Alfworld has six different environment types: 1. clean, 2. heat, 3. cool, 4. examine, 5. put, 6. puttwo.
The ‘clean’ task, e.g. Task: Put a clean apple on table, requires the agent to first find the apple, then clean it (in the sink/basin) and then put it on a table.
The ‘heat’ task, e.g. Task: Put a hot pie on table, requires the agent to first find the pie, then heat it (on the stove/burner) and then put it on a table.
The ‘cool’ task, e.g. Task: Put a cool tomato on table, requires the agent to first find the tomato, then cool it (with the fridge) and then put it on a table.
The ‘examine’ task, e.g. Task: Examine the mug with the desklamp, requires the agent to first find the mug, then find the desk lamp, and then use the desk lamp.
The ‘put’ task, e.g. Task: Find some apple and put it in sidetable, requires the agent to first find an apple, and then put it on the side table.
The ‘puttwo’ task, e.g. Task: Put two cellphone in sofa, requires the agent to first find one cellphone, and then put it on the sofa, and then to find the second one and put it on the sofa.
A.2 Action Types
Alfworld has the following valid actions: 1. go to, 2. open, 3. close, 4. put, 5. take, 6. cool, 7. heat, 8. use.
go to <place>
Example: go to table 1
open <object>
Example: open door 1
close <object>
Example: close door 1
put <object> in/on <place>
Example: put apple 1 in/on table 1
take <object> from <place>
Example: take apple 1 from table 1
cool <object> with <place>
Example: cool apple 1 with fridge 1
heat <object> with <place>
Example: heat apple 1 with fire 1
use <object>
Example: use desklamp 1
A.3 License
Alfworld has the permissible MIT license; we used it in line with the license.
Appendix B Webshop
B.1 Commands and environment
Webshop has one environment type: ‘search & buy’, as well as two commands: 1. search, 2. click.
click[<button>]
Example: click[< Back to Search]
search[<query>]
Example: search[interesting book]
B.2 Prodcuts and attributes
Webshop has over 1 million real-world products across 5 main categories (fashion, makeup, electronics, furniture and food) and 113 sub-categories.
B.3 License
Webshop has the permissible Princeton license; we used it in line with the license.
Appendix C Code snippet to call OpenAI / GPT-3.5
client = openai.OpenAI(
# Defaults to os.environ.get("OPENAI_API_KEY")
# api_key=OPENAI_KEY,
full_prompt = [{
"role": "user",
"content": prompt
chat_completion = client.chat.completions.create(
model="gpt-3.5-turbo-1106",
messages=full_prompt,
temperature=0.0,
[…]
A prompt is given in Appendix H.
Appendix D Heuristic State-tracking Explained
Heuristic state tracking is based on the idea that the state can be inferred automatically if one follows the actions of the agent and observations of the environments. Specifically, the state at time $t$ for StateAct depends on the state at time t-1 and the action $a_{t-1}$ and observation $o_{t}$. For example, if the state at time $t$ the ‘current location’ of the state is set to table 1 and the action is go to fridge 1 and the observation is successful, then the ‘current location’ can be updated to be fridge 1 automatically. This rule based ‘state-tracking’ is how the heuristics work.
Appendix E ‘Step Length Analysis’ Discussion
An alternative to calculating and comparing step length could be ‘gold solutions’ to measure optimal step length and optimality of an agent. We see two issues. Firstly, the annotation cost of creating gold solutions. Secondly, it is not clear what the gold solution should be. Concretely in Alfworld, an ‘oracle’ solution could have very few steps as it would immediately go to the location of the ‘hidden’ object, while a ‘non-oracle’ expert solution would have more steps as more locations would be searched. Thus a ‘difficulty’ measure would be needed instead, but it is ambiguous.
Appendix F Potential Future Work Directions
We found that ‘thoughts’ or explicit reasoning do not always help performance. It would be very interesting to systematise ‘thought’ and ‘states’ and to understand what contributes positively and the reasons why. Also, inspired by the positive results of StateAct, it is interesting to see what other improvements can be made without resorting to training, larger models or external tools. Finally, problems related to domain-specific syntax are also an interesting avenue for future work.
Appendix G StateAct Json Format
We translate the text-based StateAct prompt:
>goal: put a hot apple in fridge
current location: starting location
current inventory: None
thought: To solve the task, I need to find and take an apple, then heat it with microwave, then put it in fridge. First I need to find an apple. An apple is more likely to appear in fridge (1), diningtable (1), coffeetable (1), drawer (1), cabinet (1-13), garbagecan (1). I can check one by one, starting with fridge 1.
action: go to fridge 1
Into the following json format:
>{"goal": "put a hot apple in fridge",
"current_location": "starting location",
"current_inventory": None,
"thought": "To solve the task, I need to find and take an apple, then heat it with microwave, then put it in fridge. First I need to find an apple. An apple is more likely to appear in fridge (1), diningtable (1), coffeetable (1), drawer (1), cabinet (1-13), garbagecan (1). I can check one by one, starting with fridge 1."
"action": "go to fridge 1"
Appendix H StateAct Prompts
H.1 Example Alfworld prompt for the ‘heat’ task
Interact with a household to solve a task.
Here are 2 examples:
You are in the middle of a room. Looking quickly around you, you see a cabinet 13, a cabinet 12, a cabinet 11, a cabinet 10, a cabinet 9, a cabinet 8, a cabinet 7, a cabinet 6, a cabinet 5, a cabinet 4, a cabinet 3, a cabinet 2, a cabinet 1, a coffeemachine 1, a countertop 1, a diningtable 1, a drawer 1, a fridge 1, a garbagecan 1, a microwave 1, a shelf 3, a shelf 2, a shelf 1, a sinkbasin 1, a stoveburner 4, a stoveburner 3, a stoveburner 2, a stoveburner 1, and a toaster 1.
Your task is to: put a hot apple in fridge.
>goal: put a hot apple in fridge
current location: starting location
current inventory: None
thought: To solve the task, I need to find and take an apple, then heat it with microwave, then put it in fridge. First I need to find an apple. An apple is more likely to appear in fridge (1), diningtable (1), coffeetable (1), drawer (1), cabinet (1-13), garbagecan (1). I can check one by one, starting with fridge 1.
action: go to fridge 1
The fridge 1 is closed.
>goal: put a hot apple in fridge
current location: fridge 1
current inventory: None
thought: None
action: open fridge 1
You open the fridge 1. The fridge 1 is open. In it, you see a cup 1, and a egg 1.
>goal: put a hot apple in fridge
current location: fridge 1
current inventory: None
thought: None
action: go to diningtable 1
On the diningtable 1, you see a apple 1, a bread 1, a fork 3, a fork 2, a fork 1, a mug 2, a peppershaker 3, a plate 2, a pot 1, a soapbottle 3, a spatula 1, a spoon 3, a spoon 2, a spoon 1, a winebottle 3, and a winebottle 2.
>goal: put a hot apple in fridge
current location: diningtable 1
current inventory: None
thought: Now I find an apple (1). Next, I need to take it.
action: take apple 1 from diningtable 1
You pick up the apple 1 from the diningtable 1.
>goal: put a hot apple in fridge
current location: diningtable 1
current inventory: apple 1
thought: Now I take an apple (1). Next, I need to go to a microwave (1) and heat it.
action: go to microwave 1
The microwave 1 is closed.
>goal: put a hot apple in fridge
current location: microwave 1
current inventory: apple 1
thought: None
action: heat apple 1 with microwave 1
You heat the apple 1 using the microwave 1.
>goal: put a hot apple in fridge
current location: microwave 1
current inventory: apple 1
thought: Now I heat an apple (1). Next, I need to put it in/on fridge 1.
action: go to fridge 1
The fridge 1 is open. In it, you see a cup 1, and a egg 1.
>goal: put a hot apple in fridge
current location: fridge 1
current inventory: apple 1
thought: None
action: put apple 1 in/on fridge 1
You are in the middle of a room. Looking quickly around you, you see a cabinet 10, a cabinet 9, a cabinet 8, a cabinet 7, a cabinet 6, a cabinet 5, a cabinet 4, a cabinet 3, a cabinet 2, a cabinet 1, a coffeemachine 1, a countertop 3, a countertop 2, a countertop 1, a diningtable 1, a drawer 6, a drawer 5, a drawer 4, a drawer 3, a drawer 2, a drawer 1, a fridge 1, a garbagecan 1, a microwave 1, a sinkbasin 1, a stoveburner 4, a stoveburner 3, a stoveburner 2, a stoveburner 1, and a toaster 1.
Your task is to: heat some egg and put it in diningtable.
>goal: heat some egg and put it in diningtable
current location: starting location
current inventory: None
thought: To solve the task, I need to find and take an egg, then heat it with microwave, then put it in diningtable. First I need to find an egg. An egg is more likely to appear in fridge (1), countertop (1-3), diningtable (1), stoveburner (1-4), toaster (1), garbagecan (1), cabinet (1-10). I can check one by one, starting with fridge 1.
action: go to fridge 1
The fridge 1 is closed.
>goal: heat some egg and put it in diningtable
current location: fridge 1
current inventory: None
thought: None
action: open fridge 1
You open the fridge 1. The fridge 1 is open. In it, you see a lettuce 2, a mug 2, and a potato 3.
>goal: heat some egg and put it in diningtable
current location: fridge 1
current inventory: None
thought: None
action: go to countertop 1
On the countertop 1, you see a bread 1, a fork 1, and a saltshaker 1.
>goal: heat some egg and put it in diningtable
current location: countertop 1
current inventory: None
thought: None
action: go to countertop 2
On the countertop 2, you see nothing.
>goal: heat some egg and put it in diningtable
current location: countertop 2
current inventory: None
thought: None
action: go to countertop 3
On the countertop 3, you see a bowl 1, a butterknife 1, a egg 2, a kettle 2, a plate 1, a sink 1, and a spatula 2.
>goal: heat some egg and put it in diningtable
current location: countertop 3
current inventory: None
thought: Now I find an egg (2). Next, I need to take it.
action: take egg 2 from countertop 3
You pick up the egg 2 from the countertop 3.
>goal: heat some egg and put it in diningtable
current location: countertop 3
current inventory: egg 2
thought: Now I take an egg (2). Next, I need go to a microwave (1) and heat it.
action: go to microwave 1
The microwave 1 is closed.
>goal: heat some egg and put it in diningtable
current location: microwave 1
current inventory: egg 2
thought: None
action: heat egg 2 with microwave 1
You heat the egg 2 using the microwave 1.
>goal: heat some egg and put it in diningtable
current location: microwave 1
current inventory: egg 2
thought: Now I heat an egg (2). Next, I need to put it in/on diningtable 1.
action: go to diningtable 1
On the diningtable 1, you see a apple 2, a bread 3, a egg 1, a kettle 1, a knife 1, a mug 1, a papertowelroll 1, a peppershaker 2, a potato 1, a soapbottle 1, and a spatula 1.
>goal: heat some egg and put it in diningtable
current location: diningtable 1
current inventory: egg 2
thought: None
action: put egg 2 in/on diningtable 1
Here is the task.
<CURRENT TASK>
H.2 Example Webshop prompt
Webshop
Instruction:
i would like a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
[Search]
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Search Home Page
Current Selection: None
Thought: None
Action: search[3 ounce bright citrus deodorant sensitive skin]
Observation:
[Back to Search]
Page 1 (Total results: 50)
[Next >]
[B078GWRC1J]
Bright Citrus Deodorant by Earth Mama | Natural and Safe for Sensitive Skin, Pregnancy and Breastfeeding, Contains Organic Calendula 3-Ounce
\$10.99
[B078GTKVXY]
Ginger Fresh Deodorant by Earth Mama | Natural and Safe for Sensitive Skin, Pregnancy and Breastfeeding, Contains Organic Calendula 3-Ounce
\$10.99
[B08KBVJ4XN]
Barrel and Oak - Aluminum-Free Deodorant, Deodorant for Men, Essential Oil-Based Scent, 24-Hour Odor Protection, Cedar & Patchouli Blend, Gentle on Sensitive Skin (Mountain Sage, 2.7 oz, 2-Pack)
\$15.95
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Search Results Page, page 1
Current Selection: None
Thought: B078GWRC1J and B078GTKVXY are bright citrus deodorant less then 50 dollars. I can check B078GWRC1J first.
Action: click[B078GWRC1J]
Observation:
[Back to Search]
[< Prev]
scent [assorted scents][bright citrus][calming lavender][ginger fresh][simply non-scents]
size [travel set (4-pack)][3 ounce (pack of 1)][3-ounce (2-pack)]
Bright Citrus Deodorant by Earth Mama | Natural and Safe for Sensitive Skin, Pregnancy and Breastfeeding, Contains Organic Calendula 3-Ounce
Price: \$10.99
Rating: N.A.
[Description]
[Features]
[Reviews]
[Buy Now]
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Item B078GWRC1J Page
Current Selection: None
Thought: For 3 ounce bottle of bright citrus deodorant for sensitive skin, the item has options bright citrus and ’3 ounce (pack of 1)’ and seems good to buy.
Action: click[bright citrus]
Observation: You have clicked bright citrus.
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Item B078GWRC1J Page
Current Selection: bright citrus
Thought: None
Action: click[3 ounce (pack of 1)]
Observation: You have clicked 3 ounce (pack of 1).
Goal: Buy a 3 ounce bottle of bright citrus deodorant for sensitive skin, and price lower than 50.00 dollars
Current Location: Item B078GWRC1J Page
Current Selection: bright citrus, 3 ounce (pack of 1)
Thought: None
Action: click[Buy Now]